# “Rich-Get-Richer”? Analyzing Content Creator Earnings Across Large Social Media Platforms

### v1

@@ 4.1  Restatement of core findings
Our analysis confirms that earnings distributions on most of the major social media platforms closely follow power law dynamics, with an average power law exponent ($\alpha$) value around 2.0. One implication of this is that, as per the preferential attachment model of [6], initial advantage online generates further ones. Patreon content creator earnings across major social media platforms are driven by compounding mechanisms resembling capital income rather than labor-based dynamics.
@@
Temporal trends show worsening (and converging) earnings power law earnings dynamics of Patreon creators on Instagram, YouTube, Facebook, and Twitch, where decreasing $\alpha$ values suggest growing dominance of top Patreon creators. Instagram’s $\alpha$ dropped from around 2.1 in 2018 to around 1.84 in 2023, coinciding with accelerated algorithmically-driven content prioritization. These trends may indicate that algorithmic systems increasingly favor viral content over diverse creator engagement, potentially undermining long-term ecosystem health. Twitter’s power law relation weakened during the period of Musk’s takeover (2021 vs. 2024), indicating no harm to independent Patreon creators as a whole on Twitter - with NSFW creators flocking to the platform (and leaving Instagram).
Platforms with lower $\alpha$ values, such as YouTube and Instagram, face significant risks, including discouraging mid-tier creators and amplifying the spread of viral or harmful content. Conversely, platforms like Patreon and Twitter, with higher $\alpha$ values, may be able to foster healthier ecosystems, supporting broader creator participation.
4.2  Market power and earnings inequality
What is less clear is the relationship between the earnings distribution, seen within the platforms we have examined, and how competitive the social media market is – also known as the underlying ‘market structure’ (defined as the number of firms in the market, driven by cost structure and ultimately technology) [13].
If Patreon and Twitter compete less with the other social media platforms, then this may be evidence of the relationship between a less concentrated market structure and more equal algorithmic distribution of content creator earning - since Patreon only& Twitter only (Patreon content creator) earnings on these platforms is more equal. Platforms that are more dominant in the social media market may tend towards less equal distributions (YouTube vs. Instagram). But it is not necessarily obvious why promoting viral algorithmic content is necessarily more profitable to the platform.
More competition in the social media market has had ambiguous effects, theoretically, on the content creator earnings distribution. Greater competition in the social media platforms market from TikTok occurred through it promoting more longer-tail content – but this was purely algorithmically selected content (unrelated to the user’s social graph) [11]. This shifted competition directly to the algorithm, which over time seems to have promoted more addictive content, and perhaps in turn more viral content (thereby fostering a ‘winner takes all’ dynamic). The counter, egalitarian-enhancing tendency, from more competition in the market, has been that platforms have increased pay to content creators to attract them away from TikTok [25]. But this could just as much increase within platform inequality if top-tier content creators are the focus of such competitive poaching.
Ultimately, the most likely explanation may reflect technological factors more than market structure. More inequality among content creators on – and even between – platforms may simply reflect platforms ads-sharing model with content creators. By paying content creators a percentage of ads share, they are receiving a share of ads sales, that grows exponentially with sales (algorithmic attention). They are effectively sharing in algorithmic allocations directly. It is notable that YouTube offers the most generous share of ad sales and it hosts the greatest earnings inequality among Patreon earners [26].1010endnote: 10Though its unclear if this mechanism would be reflected in our dataset, since Patreon content creators monetize through clicks through to Patreon. Applied to our specific dataset, the argument similarly becomes that Patreon content creators generate capital-like inequality of earnings within, and between, their primary platforms due to content creation and distribution being an ultra-low marginal cost endeavour, thereby favoring economies of scale (and favoring network effects).
A central challenge facing recommender systems on mature platforms is their potential to compound initial advantage (viewership), by exploiting popular, existing material and focusing less on untested ‘fresh’ content [17, 60]. To the extent that this becomes caked into a platform’s business model, with associated commercial incentives, it can become difficult to dislodge this approach through market forces alone.
A “longer-tail” algorithmic system [2], whereby a more diverse range of content is explored and promoted, invariably trades off some (more) immediate profit (exploitation of existing content), in favor of greater content exploration [35, 65]. But maximizing for immediate user engagement [39], as a quarterly economic imperative facing the platform’s shareholders, can lead to less new content being created and even less longer-term user engagement – even if potentially at the expense of diminished immediate user satisfaction [45, 65]. Several potential reforms are possible to grow healthier – more diverse and equitable – ecosystems of content promotion and consumption:
Firstly, we recommend measures that incentivize long-term orientated (algorithmic) A/B testing, to help highlight the business case for longer-term business optimizations at the algorithmic level [44]. [35] recommends modifying the learning rates of recommendation algorithms to help balance the trade-off between exploration and exploitation, ensuring that high-quality content is recognized and then promoted quickly. Recent research suggests that incorporating mechanisms to surface underrepresented (“longer-tail”) content [2] could enhance overall platform performance without necessarily sacrificing user satisfaction [34, 65]. Use of longer-tail discovery algorithms on social media to unlock and match a greater variety of non-viral content to users might also be incentivized through mandatory platform mechanisms that penalize the creation of low-quality content from dominant creators [35]. Facebook has recently chosen to do the opposite [62]
Our second recommendation is for greater platform disclosure of the drivers of their recommendation systems, in-line with the Digital Services Act (DSA, Article 27) of the European Union – and in particular as a means to incentivize greater use of long-tail recommender systems [52]. Platforms should be required to disclose key details about their recommendation algorithms to regulators, researchers, and the public. These disclosures should include the algorithmic weighting of past engagement metrics, such as views and likes, which often amplify the visibility of previously successful content. In addition, platforms should clarify the criteria by which new or less popular content is surfaced within the recommendation engine, ensuring greater accountability for algorithmic bias. In addition, mechanisms in place that punish low quality, click-bait content, should be ideally detailed . In the DSA mandates platform audits, the extent to which AI and ML recommendation systems favor certain types of creators, content genres, or user demographics should be evaluated .
Finally, platforms could allow users to customize their feeds, such as through “content diversity sliders” to adjust the balance between popular and niche content. YouTube, for example, sometimes tries to encourage this by asking if you want to explore unrelated topics to what you have been watching. And although Article 27 of the DSA requires platforms to offer users such “non-profiling” alternatives (e.g. chronological feeds), this is seen as an alternative to algorithmically curated ones.
Lastly, reforms must take into account emerging AI technologies. AI technologies already impact the allocation of attention and in turn earnings within this platforms. AI drives more of the recommendations now seen online [29], boosting monetization [58]. AI-generated content has the potential to foster more egalitarian ecosystems [67], lowering barriers to entry. But more than this, as AI-generated content becomes more central to generating both organic and advertising content, attribution mechanisms to ensure copyright holders of the underlying content are rewarded will become important to fostering fairly remunerated ecosystems [48]. This also highlights the importance of connecting platform content creator ecosystems to seemingly external ones (underlying copyright holders or content creator owners).
@@ Appendix A Robustness and Further Analysis
Table 2: Further statistics of Patreon monthly earnings by platform with NSF
@@
Note: This graph compares tail exponents estimated using the Hill estimator, the Clauset-Newman-Shalizi (CNS) method [16], and the estimators described and implemented by [64]. For the latter, the adjusted Hill estimator (Adj alpha) and the moments estimation (Moments) are used. The CNS method, as introduced by [16], relies on maximum likelihood estimation to fit the power-law model and uses the Kolmogorov-Smirnov (KS) statistic to assess goodness-of-fit. In contrast, [64]’s implementation incorporates an automatic double bootstrapping procedure to determine the minimum value threshold.
@@ Appendix B Preferential Attachment Model
B.1  Derivation of the relationship between $\alpha$ and p
We provide a heuristic derivation of the power law. Based on [46, 21, 14], the derivation of the power law begins with an approximation to make the model tractable. We switch from a probabilistic to a deterministic framework, where we model the growth of links to a node as a continuous function of time, $x_{j}(t)$, for node $j$ created at time $t=j$.
The growth rate of $x_{j}(t)$ is governed by the differential equation:
$\frac{dx_{j}}{dt}=\frac{p}{t}+(1-p)\frac{x_{j}}{t}$
This equation reflects the combined effect of random linking (with probability $p$) and preferential attachment (with probability 1p1-p). Solving this differential equation yields:
$x_{j}(t)=\left(\frac{t}{j}\right)^{1-p}-1$
indicating that the number of links to a node grows as a function of the ratio of the current time to the node’s creation time, raised to the power of 1p1-p.
The key insight from solving the deterministic model is that the fraction of nodes with a given number of links $k$ at time $t$ follows a power law distribution. Specifically, the probability that a node has $k$ links decays as $k^{-\alpha}$, where $\alpha=1+\frac{1}{1-p}$, a direct consequence of the rich-get-richer mechanism.
This derivation not only highlight the origin of the power law in network connectivity but also highlights the dependency of the power law exponent on the parameter $p$, which modulates the balance between random and preferential linking. As $p$ decreases, emphasizing preferential attachment, the exponent $\alpha$ approaches 2, indicating a stronger rich-get-richer effect with a more pronounced tail in the distribution, allowing for a higher probability of nodes with very large numbers of links.
@@ B.2  Derivation of the power law exponent in the original Barabási-Albert model
The preferential attachment model [6] is based on two mechanisms: growth, where a new node is added at each time step, and preferential attachment, where each new node connects to $m$ existing nodes with a probability proportional to their degrees.
@@
Substituting back into the equation for $k_{i}(t)$, we obtain
@@
The cumulative distribution function (CDF) $P(k)$ is the probability that a node has degree greater than or equal to $k$ at time $t$. From the degree evolution equation, we have
$k_{i}(t)=m\left(\frac{t}{t_{i}}\right)^{1/2}.$ (12)
Solving for $t_{i}$, we get
@@
which shows that the CDF follows a power law with exponent $\alpha=2$.

### v2

@@ 5  Conclusion
Our analysis confirms that earnings distributions on most of the major social media platforms closely follow power law dynamics, with an average power law exponent ($\alpha$) value around 2.0. One implication of this is that, as per the preferential attachment model of [barabasi1999emergence], initial advantage online generates further ones. Patreon content creator earnings across major social media platforms are driven by compounding mechanisms resembling highly concentrated capital income rather than labor income based dynamics.
@@
Temporal trends show worsening (and converging) power law earnings dynamics of Patreon creators on Instagram, YouTube, Facebook, and Twitch, where decreasing $\alpha$ values suggest growing dominance of top Patreon creators. Instagram’s $\alpha$ dropped from around 2.1 in 2018 to around 1.84 in 2024, coinciding with accelerated algorithmically-driven content prioritization. These trends may indicate that algorithmic systems increasingly favor viral content over diverse creator engagement, potentially undermining long-term ecosystem health. Platforms with lower $\alpha$ values, such as YouTube and Instagram, face significant risks, including discouraging mid-tier creators and amplifying the spread of viral or harmful content.
It is important to restate that our findings offer only a very partial view of the actual (total) earnings distributions of content creators within social media platforms, since all major types of earnings are excluded. They are best interpreted as estimating the attention and earnings dynamics between platforms, on the basis of their algorithmic allocations of attention.
Finally, consideration of emerging AI technologies on algorithmic behavior and market structure are vital. AI technologies have for some time underpinned recommendation algorithms and advertising algorithms [covington2016deep, schafer_2023, gairola_2024]. If AI drives an explosion in content creation this risks reinforcing the dominance of existing platforms that intermediate and filter such content, as we come to rely on their algorithms for discovery even further. It may also lower barriers to entry for content creators, but it will not change online inequality and concentration in attention and earnings unless the underlying algorithmic mechanisms changes – even if AI-generated content has the potential to foster more egalitarian ecosystems [zhou2024generative]. It’s unclear at this stage how AI-native interfaces will change these dynamics.
@@ Appendix A Robustness and Further Analysis
Table 2: Further statistics of Patreon monthly earnings by platform with NSFW
@@
Note: This graph compares tail exponents estimated using the Hill estimator, the Clauset-Newman-Shalizi (CNS) method [clauset2009power], and the estimators described and implemented by [voitalov2019scale]. For the latter, the adjusted Hill estimator (Adj alpha) and the moments estimation (Moments) are used. The CNS method, as introduced by [clauset2009power], relies on maximum likelihood estimation to fit the power-law model and uses the Kolmogorov-Smirnov (KS) statistic to assess goodness-of-fit. In contrast, [voitalov2019scale]’s implementation incorporates an automatic double bootstrapping procedure to determine the minimum value threshold.
@@ Appendix B Preferential Attachment Model
B.1  Derivation of the relationship between $\alpha$ and $\gamma$
We provide a heuristic derivation of the power law. Based on [mitzenmacher2004brief, easley2010networks, chung2003generalizations], the derivation of the power law begins with an approximation to make the model tractable. We switch from a probabilistic to a deterministic framework, where we model the growth of links to a node as a continuous function of time, $x_{j}(t)$, for node $j$ created at time $t=j$.
The growth rate of $x_{j}(t)$ is described by the differential equation
$\frac{dx_{j}}{dt}=\frac{\gamma}{t}+(1-\gamma)\frac{x_{j}}{t},$
which combines two forces: a baseline flow of random links (scaled by $\gamma$) and a preferential term that is proportional to the current number of links (scaled by $1-\gamma$). This is a first-order linear differential equation, which we can rewrite as
$\frac{dx_{j}}{dt}-\frac{1-\gamma}{t}x_{j}=\frac{\gamma}{t}.$
Using the integrating factor $t^{-(1-\gamma)}$, we obtain
$\frac{d}{dt}\bigl[x_{j}(t)\,t^{-(1-\gamma)}\bigr]=\gamma\,t^{-(2-\gamma)}.$
Integrating with respect to $t$ and imposing the initial condition $x_{j}(j)=0$ (a node has no incoming links at birth) gives
$x_{j}(t)=\frac{\gamma}{1-\gamma}\left[\left(\frac{t}{j}\right)^{1-\gamma}-1\right].$
Thus, in this deterministic approximation, the expected number of links for node $j$ increases as a power of the ratio between current time $t$ and its creation time $j$, with exponent $1-\gamma$.
The key insight from solving this model is that the fraction of nodes with a given number of links $k$ at time $t$ follows a power law distribution. Specifically, the probability that a node has $k$ links decays as $k^{-\alpha}$, where $\alpha=1+\frac{1}{1-\gamma}$, a direct consequence of the rich-get-richer mechanism.
This derivation not only highlight the origin of the power law in network connectivity but also highlights the dependency of the power law exponent on the parameter $\gamma$, which controls the balance between random and preferential linking. As $\gamma$ decreases, emphasizing preferential attachment, the exponent $\alpha$ approaches 2, indicating a stronger rich-get-richer effect with a more pronounced tail in the distribution.
@@ B.2  Derivation of the power law exponent in the original Barabási-Albert model
The preferential attachment model [barabasi1999emergence] is based on two mechanisms: growth, where a new node is added at each time step, and preferential attachment, where each new node connects to $m$ existing nodes with a probability proportional to their degrees.
@@
Substituting back into the equation for $k_{i}(t)$, we obtain the degree evolution equation:
@@
From the degree evolution equation, solving for $t_{i}$, we get
@@
which shows that the CDF follows a power law with exponent of 2. Differentiating this tail distribution shows that the degree density satisfies
$p(k)\propto k^{-3},$ (15)
so in the notation used in the main text, where we write $p(k)\propto k^{-\alpha}$, the Barabási–Albert model corresponds to $\alpha=3$.
Appendix C Missing Data and Imputation Validation
This appendix quantifies the extent of missing earnings data and evaluates the accuracy of the imputation method used in the analysis. Roughly one-third of creator-level earnings are missing. On Patreon in particular, creators can choose not to display their earnings publicly, and this choice is systematically related to their scale: higher-earning creators with more paid subscribers are disproportionately likely to hide their earnings. If we were to drop these observations, the right tail of the earnings distribution would be mechanically truncated, leading to a downward bias in both estimated inequality and the fitted power-law exponent.
To mitigate this, we impute missing earnings using an OLS model estimated on creators with observed earnings:
$\texttt{Earnings}\sim\texttt{Paid.Members}+\texttt{All.Members}+\texttt{Is.Nsfw}+\texttt{Year}+\texttt{Category},$
fit on complete cases (na.omit). The regressors capture the main observable drivers of earnings: the scale of the paying audience (Paid.Members), the broader follower base (All.Members), content type and NSFW status, and time effects. Since the probability of non-disclosure is strongly related to these observables (especially paid members), this specification approximates a setting in which earnings are missing at random conditional on the covariates. The fitted model is then used only to predict missing earnings; observed earnings remain unchanged.
Table 3 reports the sample size and missingness, alongside model performance. We show in-sample fit ($R^{2}$) and error magnitudes in levels. For interpretability, each error metric also reports its cross-validated counterpart in parentheses, computed via 10-fold out-of-sample prediction using the same specification. The model explains a substantial share of observed variation ($R^{2}\!=\!0.78$). Errors are moderate in absolute terms (RMSE $\approx\mathdollar 547$, MAE $\approx\mathdollar 81$), and the cross-validated errors are very similar, indicating limited overfitting. Given that our power-law estimates are driven by the broad shape of the upper tail rather than exact prediction of individual creators, this level of accuracy is sufficient for our purposes and preferable to discarding systematically missing high-earning observations.
Table 3: Missing data and imputation diagnostics
N_Tot N_Obs N_Miss Miss (%) $R^{2}$ RMSE (CV) MAE (CV)
595,912 390,458 205,454 34.5% 0.778 547.17 (530.55) 80.89 (81.17)
Notes. N_Tot: total creators before imputation; N_Obs: non-missing earnings; N_Miss: missing earnings; Miss (%): share missing. Values in parentheses are 10-fold cross-validated (out-of-sample) metrics. RMSE = root mean squared error; MAE = mean absolute error; absolute-error metrics are in USD.